OpenStudy (anonymous):

Simplify: 2(sqrt -24 )- (sqrt -54 )

OpenStudy (anonymous):

havent u learned it yet

OpenStudy (anonymous):

can you explain with out giving me an answer please

OpenStudy (anonymous):

sqrt of 24 is whaT?

OpenStudy (anonymous):

That's -24, right?

OpenStudy (anonymous):

Let's assume it is and use i to get rid of the minus signs (if they are really positive just lose the i after). And while we are at it, lets split the roots up into smaller ones to get \[i (2\sqrt{6}\sqrt{4} - \sqrt{6}\sqrt{9})\] You can do it now, right?

jhonyy9 (jhonyy9):

radical -24 = 2i radical 6 radical -54 = 3i radical 6 2(2i radical 6 ) - ( 3i radical 6 ) = 2 (i radical 6 ( 2-3 ) = -2i radical 6

OpenStudy (anonymous):

Um, it's just -i root6 isn't it?

OpenStudy (anonymous):

I mean i root 6